# Identification of Dynamic Nonlinear Panel Models

### v2

@@ A.10 Dynamic Censored Models with Latent Dependent Variables
Consider the following dynamic censored models with the latent lagged outcome $Y^{*}_{i,t-1}$:
$Y^{*}_{it}$ $=Z_{it}^{\prime}\beta_{0}+Y^{*}_{i,t-1}\gamma_{0}+\alpha_{i}+\epsilon_{it},$
$Y_{it}$ $=\max\{Y^{*}_{it},0\},$
@@
In this model, the endogenous variable $X_{it}$ is the lagged outcome: $X_{it}=Y^{*}_{i,t-1}$. However, the variable $Y^{*}_{i,t-1}$ is not observed in data, so the results in Proposition 8 cannot be directly applied here. Due to this feature in the dynamic model, we need to adjust the results in Proposition 8. Given that $Y^{*}_{i,t-1}=Y_{i,t-1}$ when $Y_{i,t-1}>0$, we can further relax the lower and upper bounds in (8) to identify $\theta_{0}$.
@@
$\Pr(Y_{it}\leq z_{t}^{\prime}\beta+Y^{*}_{i,t-1}\gamma-a\mid z_{st})\\ \geq\Pr(Y_{it}\leq z_{t}^{\prime}\beta+Y_{i,t-1}\gamma-a,Y_{i,t-1}>0\mid z_{st}):=L_{t,cen}(a\mid z_{st};\theta).$
@@
$\Pr(0<Y_{is}\leq z_{s}^{\prime}\beta+Y^{*}_{i,s-1}\gamma-a\mid z_{st})+\Pr(Y_{is}=0\mid z_{st})\leq U_{s,cen}(a\mid z_{st};\theta),$
where $U_{s,cen}(a\mid z_{st};\theta)$ is defined as
$U_{s,cen}(a\mid z_{st};\theta):=\Pr(0<Y_{is}\leq z_{s}^{\prime}\beta+Y_{i,s-1}\gamma-a,Y_{i,s-1}>0\mid z_{st})\\ +\Pr(Y_{is}>0,Y_{i,s-1}=0\mid z_{st})+\Pr(Y_{is}=0\mid z_{st}).$
@@
Under Assumption 1 and $X_{it}=Y_{i,t-1}^{*}$, the identified set for $\theta_{0}$ is characterized by the set of parameters $\theta$ that satisfy the following conditions:
$U_{s,cen}(a\mid z_{st};\theta)\geq L_{t,cen}(a\mid z_{st};\theta).$
for any $a\in\mathcal{R}$, $s,t\leq T$, and $z_{st}$.

### v3

@@ A.8 Dynamic Censored Models with Latent Dependent Variables
Consider the following dynamic censored models with the latent lagged outcome $Y_{i,t-1}^{*}$:
$Y_{it}^{*}$ $=Z_{it}^{\prime}\beta_{0}+Y_{i,t-1}^{*}\gamma_{0}+\alpha_{i}+\epsilon_{it},$
$Y_{it}$ $=\max\{Y_{it}^{*},0\},$
@@
In this model, the endogenous variable $X_{it}$ is the lagged outcome: $X_{it}=Y_{i,t-1}^{*}$. However, the variable $Y_{i,t-1}^{*}$ is not observed in data, so the results in Proposition 9 cannot be directly applied here. Due to this feature in the dynamic model, we need to adjust the results in Proposition 9. Given that $Y_{i,t-1}^{*}=Y_{i,t-1}$ when $Y_{i,t-1}>0$, we can further relax the lower and upper bounds in (9) to identify $\theta_{0}$.
@@
$\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta+Y_{i,t-1}^{*}\gamma-c\mid z)\\ \geq\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta+Y_{i,t-1}\gamma-c,Y_{i,t-1}>0\mid z):=L_{t,cen}(c\mid z;\theta).$
@@
$\mathbb{P}(0<Y_{is}\leq z_{s}^{\prime}\beta+Y_{i,s-1}^{*}\gamma-c\mid z)+\mathbb{P}(Y_{is}=0\mid z)\leq U_{s,cen}(c\mid z;\theta),$
where $U_{s,cen}(c\mid z;\theta)$ is defined as
$U_{s,cen}(c\mid z;\theta):=\mathbb{P}(0<Y_{is}\leq z_{s}^{\prime}\beta+Y_{i,s-1}\gamma-c,\ Y_{i,s-1}>0\mid z)\\ +\mathbb{P}(Y_{is}>0,\ Y_{i,s-1}=0\mid z)+\mathbb{P}(Y_{is}=0\mid z).$
@@
Under Assumption 1 and $X_{it}=Y_{i,t-1}^{*}$, $\theta_{0}\in\tilde{\Theta}_{I,cen}$, where the identified set $\tilde{\Theta}_{I,cen}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ such that
$\max_{t=1,...,T}L_{t,cen}(c\mid z;\theta)\leq\min_{s=1,...,T}U_{s,cen}(c\mid z;\theta).$
for any $c\in\mathcal{R}$ and any realization $z=(z_{1},...,z_{T})$ in the support of $Z_{i}$.